Siver Manufacturing paid $75,000 (no residual value) for equipment with a useful life of 5 years on January 1.  Silver uses the double-declining-balance method of depreciation. However, after two years, Silver changes to the straight-line method for depreciation purposes. What is depreciation expense in year 3?    

Select one:

a. $5,400

b. $15,000

c. $13,500

d. $9,000

e. $16,000
Answer #1

Double decline rate = 100/5*2 = 40%

Book value after two year = 75000*60%*60% = 27000

Depreciation expense 3 year = 27000/3 = 9000

So answer is d) $9000